# SLAM&Render: A Benchmark for the Intersection Between Neural Rendering, Gaussian Splatting and SLAM https://samuel-cerezo.github.io/SLAM&Render

### v2

@@ IV-B Dataset structure
  • rgb/: folder containing the timestamped color images (PNG format, 3 channels, 8 bits per channel);
  • depth/: folder containing the timestamped depth images (PNG format, 1 channel, 16 bits per channel, distance scaled by a factor of 1000). The depth images are already aligned with the RGB sensor;
  • associations.txt: file listing all the temporal associations between the RGB and depth images (format: rgb_timestamp rgb_file depth_timestamp depth_file);
  • imu.txt: file containing timestamped accelerometer and gyroscope readings (format: timestamp ax ay az gx gy gz);
  • groundtruth.txt: file containing the timestamped ground-truth of the camera poses w.r.t world reference frame, expressed as a sequence of quaternions and translation vectors (format: timestamp qx qy qz qw tx ty tz) (see Sec. IV-D);
  • groundtruth_raw.csv: file provided by the MCS containing ground-truth raw data of flange marker ring trajectory;
  • extrinsics.yaml: file containing the transformation matrices of the reference frames involved in the data collection process (see Sec. IV-C);
  • instrinsics.yaml: file containing the intrinsic calibration parameters of the camera (see Sec. IV-E);
  • robot_data/flange_poses.txt: file containing the timestamped poses of the robot flange. Each pose is expressed w.r.t. the robot base as a quaternion and translation vector (format: timestamp qx qy qz qw tx ty tz);
  • robot_data/joint_positions.txt: file reporting the timestamped angles of each robot joint (format: timestamp q1 q2 q3 q4 q5 q6).
The accelerometer data are provided in $m/s^{2}$, the gyroscope data in $rad/s$, and the ground-truth and flange positions in $mm$. All sequences are also available as ROS2 bag files.
IV-C Reference frames
@@
IV-D Motion Capture System (MCS)
[…]
Figure 5: Validation of time-synchronized data comparing the angular velocity estimated from the MCS and from the IMU gyroscope measurements.
Synchronization. The data from the MCS are initially unsynchronized. To ensure temporal alignment, for each sequence we retrieved the capture starting time provided by OptiTrack. Consequently, the ground-truth file we provide includes the corresponding Unix Epoch timestamp for each pose, calculated using the corresponding capture starting time. Following the approach presented in [25], the synchronization was validated by comparing the angular velocity estimated from the MCS with the gyroscope measurements from the IMU using cross-correlation (see Fig. 5). The maximum at zero offset confirms correct alignment.
IV-E Camera calibration
To estimate the hand-eye transformation (i.e., the transformation from the robot flange to the camera projection plane), more than 200 images containing the calibration plate were captured, and the corresponding robot pose for each image was saved. Subsequently, the calibration method presented in [26] was applied. The calibration plate (a chessboard with squares of $30\times 30~{}mm$) was detected by applying a custom detection method (see Fig. 6). The intrinsics of the RealSense D435i camera were refined from the factory calibration ones.
Figure 6: Corners and fiducial markers detected on the calibration plate.
For a more accurate calibration, not only the extrinsic parameters (hand-eye transformation) were estimated, but also the intrinsic parameters. Our experiments have proven that the use of the estimated intrinsic parameters resulted in more accurate extrinsic parameters. The intrinsic parameters of the camera (focal length (fx/fy), principal point (cx/cy) and camera distortion coefficients) are shown in Table II. The extrinsic parameters are shown in the Table III.
Parameters fx fy cx cy distortion coefficients
@@ IV-E Camera calibration
TABLE II: Intrinsic parameters of the camera in [$px$] used in our dataset.
@@
TABLE III: Extrinsic parameters (hand-eye transformation) of the camera, represented as a translation vector in [$mm$] and a unit quaternion for rotation.
@@
[$px$] Trans. [$mm$] Rot. [$deg.$] Rel. transf. err. [$a.u.$]
@@
TABLE IV: Mean, standard deviation and standard error of the calibration metrics. The last two rows show the results using all images with the estimated and factory intrinsic parameters.
To compare the accuracy of the calibration using both factory and estimated intrinsic parameters, the set of calibration images was divided into several batches (10 batches with 20 images in each), and the same calibration process was applied to each batch. The accuracy of the calibration was assessed using the metrics proposed in [27], which include translation error in milimeters (Trans. [$mm$]), rotation error in degrees (Rot. [$deg.$]), and unitless relative transformation error (Rel. transf. error [$a.u.$]).
Table IV shows the mean, standard deviation, and standard error of the calibration metrics. The accuracy of the hand-eye transformation, calculated using the estimated intrinsic parameters, showed more precise results, with reduced reprojection, translation, rotation, and relative transformation errors.
V Experiments
V-A Metrics
In this work, we report the Root Mean Square Error of the Absolute Trajectory Error (ATE RMSE $\downarrow$) to assess the camera tracking accuracy. To evaluate novel view synthesis, we follow the evaluation protocol presented in [8] and report standard photometric rendering metrics such as Peak Signal-to-Noise Ratio (PSNR $\uparrow$), Structural Similarity Index Measure (SSIM $\uparrow$), and Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$). PSNR captures pixel-level differences, SSIM reflects perceptual similarity in terms of structure, luminance, and contrast [28]. LPIPS evaluates visual fidelity based on deep feature representations [29].
V-B Results
To evaluate the dataset’s suitability for benchmarking methods at the intersection of SLAM and novel view synthesis, we present results from state-of-the-art techniques.
Novel View Synthesis. To highlight the usefulness of independent test camera trajectories, we evaluated the performance of two state-of-the-art baselines: Gaussian Splatting [8] and FeatSplat [30] (this last one specifically designed to reduce viewpoint overfitting). In this experiment, setup-1, setup-3, and setup-4 were used under natural, warm, and dark lighting conditions. Baselines were initialized with random point-clouds and a black background, using 4 Spherical Harmonics bands for Gaussian Splatting, and 32-dimensional latent vectors and 1 hidden layer of 64 neurons MLP for FeatSplat. They were optimized for a total of 21k iterations, and we report metrics both at 7k and 21k iterations. The baselines were optimized on the respective train trajectories of each setup, skipping one every ten frames. We evaluate both with the common approach of using skipped frames from the train trajectory [4, 8] (see Table V) and using the independently recorded test trajectory of each setup (see Table VI).
@@ V-B Results
TABLE V: Classic evaluation of Gaussian Splatting (gsplat) and FeatSplat (fsplat) reporting PSNR[$\uparrow$], LPIPS[$\downarrow$], and SSIM[$\uparrow$].
@@
TABLE VI: Evaluation on independent test trajectory of Gaussian Splatting (gsplat) and FeatSplat (fsplat) reporting PSNR[$\uparrow$], LPIPS[$\downarrow$], and SSIM[$\uparrow$]
Figure 7: Comparison of Novel View Synthesis for validation and test viewpoints. Despite both models perform well when evaluated on validation viewpoints of the training trajectory, the independent test trajectory reveals overfitting.
The common evaluation indicates an improvement in rendering performance, for both baselines, through the complete optimization process (see Table V). Both baselines achieve their best performance under dark lighting conditions, which is probably due to using black as initial background color. Nevertheless, the test trajectory at SLAM&Render reveals that both baselines overfit to the training trajectory. Note how the rendering quality does not improve beyond the 7k iteration and even slightly declines in Table Table VI. Qualitative analysis in Fig. 7 also showcases the contrast between the high quality rendering of validation viewpoints and the artifacts generated in test viewpoints. This result validates the need for datasets with independent trajectories to evaluate novel view synthesis methods and shows that the classic approach of sampling evaluation frames from the training trajectory does not inform about trajectory overfitting.
@@
TABLE VII: ATE RMSE [$\downarrow cm$] on different sequences running three algorithm configurations: MonoGS, Kinematics and MonoGS with Kinematic data as seed
Camera tracking. To illustrate the advantage on using kinematic data, three approaches based on Gaussian Splatting SLAM [9] (MonoGS) were evaluated. In the first experiment, we used the default configuration for MonoGS. In the second experiment we incorporated the robot kinematic data directly as the camera motion. This approach is referred to as the Kinematics experiment. In the third experiment, robot kinematrics were incorporated as an initial seed for camera pose estimation within MonoGS. This is referred to as Kinem. seed experiment. Following the evaluation conducted by the authors of MonoGS in [9], the ATE RMSE for keyframes is reported in Table VII. For this experiment, setup-1, setup-3, and setup-4 were used under different light conditions while performing train and test trajectories. In all cases, it is evident that incorporating kinematic data improves camera tracking performance. However, the results also show that the naive approach of using robot kinematics as the initial seed of MonoGS decreases its accuracy. This result motivates the use of our SLAM&Render dataset for research on multimodal fusion within NeRF and Gaussian Splatting SLAM. Fig. 8 illustrates the camera tracking results for setup-4 under cold lighting condition.
@@
The trajectory using MonoGS is shown in Fig. 8a, while Fig. 8b shows the result using kinematic data as initial seed. Incorporating kinematic data allows drift correction in camera tracking.
VI Conclusions
In this work we introduced SLAM&Render, a dataset for benchmarking methods that combine SLAM and novel view rendering. Our data includes 40 sequences of synchronized RGB and depth images, IMU readings, robot kinematic data, intrinsic and extrinsic camera parameters, and ground-truth pose data across diverse scenes and conditions. SLAM&Render addresses research challenges such as generalization across trajectories and robustness to scene and illumination changes. By releasing robot kinematic data, it also enables the assessment of novel SLAM strategies when applied to robot manipulators. It aims to drive future research in learning-based and hybrid methods requiring accurate localization and multi-sensor data fusion, ultimately enhancing the deployment of systems that combine perception, mapping, and rendering.
VII Acknowledgments

### v4

@@ 4.2 Dataset structure
  • rgb/: timestamped color images (PNG format, 3 channels, 8 bits per channel);
  • depth/: timestamped depth images (PNG format, 1 channel, 16 bits per channel, distance scaled by a factor of 1000). The depth images are already aligned with the RGB sensor;
  • associations.txt: list of all the temporal associations between the RGB and depth images (format: rgb_timestamp rgb_file depth_timestamp depth_file);
  • imu.txt: timestamped accelerometer and gyroscope readings (format: timestamp ax ay az gx gy gz);
  • groundtruth.txt: timestamped ground truth of the camera poses w.r.t world reference frame, expressed as a sequence of quaternions and translation vectors (format: timestamp qx qy qz qw tx ty tz);
  • groundtruth_raw.csv: MCS ground-truth raw data of flange marker ring trajectory;
  • extrinsics.yaml: transformation matrices of the reference frames involved in the data collection process (see Sec. 4.3);
  • instrinsics.yaml: intrinsic calibration parameters of the camera;
  • robot_data/flange_poses.txt: timestamped poses of the robot flange. Each pose is expressed w.r.t. the robot base as a quaternion and translation vector (format: timestamp qx qy qz qw tx ty tz);
  • robot_data/joint_positions.txt: timestamped angles of each robot joint (format: timestamp q1 q2 q3 q4 q5 q6).
The accelerometer data are provided in $\mathrm{m/s^{2}}$, the gyroscope data in $\mathrm{rad/s}$, and the ground truth and flange positions in $\mathrm{mm}$. All sequences are also available as ROS2 bag files.
4.3 Reference frames
@@
4.4 Motion Capture System (MCS)
Calibration. To collect ground-truth data, twelve Optitrack PrimeX 13W cameras were used. These cameras were carefully positioned to focus on the robot flange from various angles, minimizing occlusion caused by the robot’s movements. The cameras were calibrated using the Motive software [^5] https://optitrack.com/software/motive/ provided by Optitrack. This process involves waving a calibration wand, throughout the area to be captured. After the wanding process, Motive compares the 2D results from the individual cameras to reconstruct a 3D capture volume. The calibration process resulted in a 3D error mean of $0.4$ mm, representing the convergence error between the twelve cameras.
Figure 5: Validation of time-synchronized data by comparing angular velocity estimates from the MCS and the IMU gyroscope measurements.
Synchronization. The data gathered from the MCS are initially unsynchronized. To ensure temporal alignment, for each sequence we retrieved the capture starting time provided by OptiTrack. It was used to compute the Unix Epoch timestamp for each pose reported in the provided ground-truth file. Following the approach presented in Bonarini et al. (2006), the synchronization was validated by comparing the angular velocity estimated from the MCS with the gyroscope measurements from the IMU using cross-correlation (see Fig. 5). The maximum at zero offset confirms correct alignment.
4.5 Camera calibration
To estimate the hand-eye transformation (i.e., the transformation from the robot flange to the camera projection plane), more than 200 images containing the calibration plate were captured, and the corresponding robot pose for each image was saved. Subsequently, the calibration method presented in Shiu and Ahmad (1989) was applied. The calibration plate (a chessboard with squares of $30\times 30~\mathrm{mm}$) was detected by applying a custom detection method (see Fig. 6). For a more accurate calibration, not only the extrinsic parameters (hand-eye transformation) were estimated, but also the intrinsic parameters. Our experiments have proven that the use of the estimated intrinsic parameters resulted in more accurate extrinsic parameters. The intrinsic parameters of the camera (focal length, principal point and camera distortion coefficients) are shown in Table 2. The extrinsic parameters are shown in the Table 3.
Figure 6: Corners and fiducial markers detected within the calibration procedure.
Table 2: Intrinsic camera parameters used in our dataset. The parameters include focal lengths (fx,fy), principal point (cx,cy), and the distortion coefficients.
Parameters fx $\mathrm{px}$] fy $\mathrm{px}$] cx $\mathrm{px}$] cy $\mathrm{px}$] distortion coefficients
@@ 4.5 Camera calibration
Table 3: Extrinsic parameters (hand–eye transformation): translation $\mathrm{mm}$] and unit quaternion rotation.
@@
Table 4: Mean, standard deviation and standard error of the calibration metrics. The last two rows show the results using all images with the estimated and factory intrinsic parameters.
@@
$\mathrm{px}$] Trans. $\mathrm{mm}$] Rot. $\mathrm{deg.}$] Rel. transf. err. $\mathrm{a.u.}$]
@@
To compare the accuracy of the calibration using both factory and estimated intrinsic parameters, the set of calibration images was divided into several batches (10 batches with 20 images in each), and the same calibration process was applied to each batch. The accuracy of the calibration was assessed using the metrics proposed in Enebuse et al. (2022), which include translation error in millimeters (Trans. [mm]), rotation error in degrees (Rot. [deg.]), and unitless relative transformation error (Rel. transf. error [a.u.]). Table 4 shows the mean, standard deviation, and standard error of the calibration metrics. The accuracy of the hand-eye transformation, calculated using the estimated intrinsic parameters, showed more accurate results, with reduced reprojection, translation, rotation, and relative transformation errors.
5 Experiments
To evaluate the dataset’s suitability for benchmarking methods at the intersection of SLAM and Novel View Synthesis, we present results from several baselines.
5.1 Metrics
We assess camera tracking performance using the Root Mean Square Error of the Absolute Trajectory Error (ATE RMSE $\downarrow$). Following Kerbl et al. (2023) to evaluate NVS, we report Peak Signal-to-Noise Ratio (PSNR $\uparrow$), Structural Similarity Index Measure (SSIM $\uparrow$), and Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$). PSNR captures pixel-level differences, while SSIM reflects perceptual similarity in structure, luminance, and contrast Wang et al. (2004). Finally, LPIPS measures perceptual fidelity using deep feature distances Zhang et al. (2018).
5.2 Results
Novel View Synthesis (NVS). To highlight the value of independent test trajectories, we evaluated two state-of-the-art baselines: Gaussian Splatting (GSPLAT) Kerbl et al. (2023) and FeatSplat (FSPLAT) Martins and Civera (2024), the latter designed to mitigate viewpoint overfitting. In this experiment, setup-1, setup-3, and setup-4 were used under natural, warm, and dark lighting conditions. Baselines were initialized with random point clouds and a black background. GSPLAT used four spherical-harmonics bands, and FSPLAT used 32-dimensional latent vectors and a multilayer perceptron with one hidden layer of 64 units. They were optimized for a total of 21k iterations, and we report metrics both at 7k and 21k iterations. The baselines were optimized on the respective train trajectories of each setup, skipping one every ten frames. We evaluate both with the common approach of using skipped frames from the train trajectory Mildenhall et al. (2021); Kerbl et al. (2023) (see Table 5) and using the independently recorded test trajectory of each setup (see Table 6).
Table 5: Classic evaluation of Gaussian Splatting (gsplat) and FeatSplat (fsplat) reporting PSNR $\uparrow$], LPIPS $\downarrow$], and SSIM $\uparrow$].
@@ 5.2 Results
Table 6: Evaluation on independent test trajectory of Gaussian Splatting (gsplat) and FeatSplat (fsplat) reporting PSNR $\uparrow$], LPIPS $\downarrow$], and SSIM $\uparrow$].
@@
The common evaluation indicates an improvement in rendering performance, for both baselines, through the complete optimization process (see Table 5). Both baselines achieve their best performance under dark lighting conditions, which is probably due to using black as initial background color. Nevertheless, the test trajectory at SLAM&Render reveals that both baselines overfit to the training trajectory. Note how the rendering quality does not improve beyond the 7k iteration and even slightly declines in Table 6. Qualitative analysis in Fig. 7 also showcases the contrast between the high quality rendering of validation viewpoints and the artifacts generated in test viewpoints. This result validates the need for datasets with independent trajectories to evaluate novel view synthesis methods and shows that the classic approach of sampling evaluation frames from the training trajectory does not inform about trajectory overfitting.
Figure 7: Comparison of NVS for validation and test viewpoints. Although both models perform well when evaluated on validation viewpoints of the training trajectory, the independent test trajectory reveals overfitting.
Table 7: ATE RMSE $\downarrow\mathrm{cm}$] across multiple sequences using three algorithm configurations: MonoGS, Kinematics and MonoGS with Kinematic data as seed.
@@
Camera tracking. To illustrate the advantage on using kinematic data, three approaches based on Gaussian Splatting SLAM Matsuki et al. (2024) (MonoGS) were evaluated. In the first experiment, we used the default configuration for MonoGS, while in the second one we integrated the robot kinematic data to describe the camera motion. The evaluation of the second approach is referred to as the Kinematics experiment. In the third experiment, robot kinematics were incorporated as an initial seed for camera pose estimation within MonoGS. This is referred to as Kinem. seed experiment. Following the evaluation protocol in Matsuki et al. (2024), the ATE RMSE for keyframes is reported in Table 7. For this experiment, setup-1, setup-3, and setup-4 were used under different light conditions while performing train and test trajectories. In all cases, it is evident that incorporating kinematic data improves camera tracking performance allowing for drift corrections. However, the results also show that the naive approach of using robot kinematics as the initial seed of MonoGS decreases its accuracy. This result motivates the use of our SLAM&Render dataset for research on multimodal fusion within NeRF and Gaussian Splatting SLAM. Fig. 8 illustrates the camera tracking results for setup-4 under cold lighting condition. The trajectory using MonoGS is shown in Fig. 8a, while Fig. 8b shows the result using kinematic data as initial seed.
@@
6 Conclusions
In this work, we introduced SLAM&Render, a dataset for benchmarking methods that combine SLAM with Neural Rendering and Gaussian Splatting. Our dataset includes 40 sequences of synchronized RGB and depth images, IMU readings, robot kinematic data, intrinsic and extrinsic camera parameters, and ground-truth pose data across diverse scenes and conditions. SLAM&Render addresses research challenges such as generalization across reproducible trajectories and robustness to scene and illumination changes. By releasing robot kinematic data, it also enables the assessment of novel SLAM strategies when applied to robot manipulators. It aims to drive future research in learning-based and hybrid methods requiring accurate localization and multi-sensor data fusion, ultimately enhancing the deployment of systems that combine perception, mapping, and rendering. Future extensions of the dataset may include dynamic scenes and a wider range of object categories, with priorities shaped by community feedback and usage patterns. These enhancements will strengthen robustness and generalization tests while maintaining the dataset’s core focus.
7 Data access methods
The datasets can be downloaded from https://samuel-cerezo.github.io/SLAM&Render. Furthermore, some example tools to load and plot the data in Python are provided and can be used as a starting point for implementations in other programming languages.
8 Acknowledgments